# Multi-Head RAG: Solving Multi-Aspect Problems with LLMs

### v1

@@ Abstract
Retrieval Augmented Generation (RAG) enhances the abilities of Large Language Models (LLMs) by enabling the retrieval of documents into the LLM context to provide more accurate and relevant responses. Existing RAG solutions do not focus on queries that may require fetching multiple documents with substantially different contents. Such queries occur frequently, but are challenging because the embeddings of these documents may be distant in the embedding space, making it hard to retrieve them all. This paper introduces Multi-Head RAG (MRAG), a novel scheme designed to address this gap with a simple yet powerful idea: leveraging activations of Transformer’s multi-head attention layer, instead of the decoder layer, as keys for fetching multi-aspect documents. The driving motivation is that different attention heads can learn to capture different data aspects. Harnessing the corresponding activations results in embeddings that represent various facets of data items and queries, improving the retrieval accuracy for complex queries. We provide an evaluation methodology and metrics, synthetic datasets, and real-world use cases to demonstrate MRAG’s effectiveness, showing improvements of up to 20% in relevance over standard RAG baselines. MRAG can be seamlessly integrated with existing RAG frameworks and benchmarking tools like RAGAS as well as different classes of data stores.
@@ 1 Introduction
Large Language Models (LLMs) transformed many machine learning tasks using in-context learning abilities. They achieved such accuracy by leveraging an increasing number of parameters, which in recent models have grown to hundreds of billions, making LLM training expensive in terms of both time and resources. It also comes with the danger of leaking confidential data into model weights [41, 34, 29]. Additionally, continuous training through fine-tuning is necessary to keep LLMs up-to-date. Even using the newest data, LLMs display an ongoing problem of hallucinations [45, 39, 14] by providing factually incorrect information. Retrieval Augmented Generation (RAG) was proposed [19, 12] in order to address these issues as well as others and make LLMs more trustworthy.
The key idea behind RAG is to enhance the generative model’s capabilities by integrating a retrieval system that can fetch relevant documents or passages from a large corpus of data. In this setting, when a query is received, the retrieval system first identifies and retrieves pertinent information, which is fed into the generative model’s context for a more accurate and relevant response. Instead of the model storing information within its weights, RAG effectively leverages external knowledge, reducing hallucinations (by grounding the LLM reply in reliable sources), and ensuring that responses contain up-to-date knowledge (e.g., by accessing the Internet), all without requiring expensive training.
@@
Intense recent research efforts have been put into RAG [11, 46, 13, 15, 42, 26, 21]. On one hand, different RAG designs have been proposed, for example RAPTOR [32], Self-RAG [3], Chain-of-Note [43], and many others [2, 7, 8, 24, 44, 36, 40]. In general, these schemes focus on making the retrieved data more accurate and relevant to the query. On the other hand, there have also been efforts into benchmarking and datasets for RAG evaluation [5, 37, 22, 9].
@@
In this work, we propose Multi-Head RAG (MRAG): a scheme that addresses the above problem. Common practice in modern RAG designs is the use of embeddings based on last-layer decoder block activations. Our key idea is to use instead the activations of the multi-head attention part of the decoder block as embeddings. The Transformer architecture can be seen as a pipeline with many (e.g., 96 for GPT-3 [6]) blocks, where a single block consists of an attention module and a feed-forward module. Each individual attention module is multi-headed: it consists of multiple parts called heads that learn different sets of weight matrices; see Figure 1 for an overview. It is conjectured that these different heads could capture different aspects of the processed data. We use this as a driving design feature that facilitates capturing the potential multi-aspectuality of the data without increasing space requirements compared to standard RAG (contribution 1).
Such multi-aspect embeddings are then directly used for both data items and query representation. Considering multi-aspectuality explicitly comes with challenges. For example, how to assess the effectiveness of a RAG solution in retrieving data that indeed does cover multiple aspects of a given domain. For this, we establish an evaluation methodology as well as a full data construction and query processing pipeline that implements the multi-aspect embedding idea (contribution 2). Our datasets facilitate broad evaluation by considering both fully-automatically generated, synthetic data and analyzing specific industry use cases that show the benefits of MRAG (contribution 3). Our evaluation illustrates the benefits in the relevance of retrieved documents, for example 20% over a modern RAG baseline for fetching multi-aspect Wikipedia articles (contribution 4). We also show how MRAG and its benchmarking principles can be seamlessly integrated with both existing RAG solutions and benchmarking frameworks such as RAGAS (contribution 5). MRAG’s code is publicly available [^1] https://github.com/spcl/MRAG.
@@ 2 The MRAG Formulation & Pipeline
Multi-Head RAG Formulation The key idea behind MRAG is simple: instead of the single activation vector generated by the last feed-forward decoder layer for the last token, we harness the $H$ separate activation vectors generated by the last attention layer for the last token, before processing it via $\mathbf{W_{o}}$. This can be formulated as a set of embeddings $\mathcal{S}=\{\mathbf{e}_{k}\forall_{k}\}$ where $\mathbf{e}_{k}=\text{head}^{k}(\mathbf{x}_{n})$, which is simply the set of all outputs from the attention heads on the last token $\mathbf{x}_{n}$ of the input. As processing with multiple heads does not change the size of the output vector, $\mathcal{S}$ has the same space requirements as standard RAG. However, because we capture the separate embeddings before their mixing with $\mathbf{W}_{o}$, we conjecture that it gives more information about what the different parts of the input attend to, facilitating capturing multi-aspectuality.
@@ 2.1.1 Data Preparation
During data preparation [Uncaptioned image], we populate a data store [Uncaptioned image] with multi-aspect MRAG text embeddings [Uncaptioned image] and their corresponding documents [Uncaptioned image] or text chunks [Uncaptioned image] (MRAG is orthogonal to the type of data being embedded, and while we primarily use chunking of documents in order to reflect modern RAG pipelines, one can also embed whole documents or even other types of data). We create the multi-aspect embedding [Uncaptioned image] of each text chunk [Uncaptioned image] using a selected decoder-based embedding model [Uncaptioned image] (this part is detailed in Section 2.2). The user of the pipeline can plug in their model [Uncaptioned image] of choice as well as use their input data. We also offer a dedicated synthetic data generator [Uncaptioned image] that can be used to construct multi-aspect input documents [Uncaptioned image] (we detail this part in Section 3) for evaluation purposes.
@@ 2.1.2 Query Execution
During query execution [Uncaptioned image], we first generate a multi-aspect embedding [Uncaptioned image] of the input query [Uncaptioned image], using the selected embedding model [Uncaptioned image] (details in Section 2.2). Then, we find the nearest multi-aspect embeddings [Uncaptioned image] and their corresponding text chunks [Uncaptioned image] in the data store [Uncaptioned image] using a special multi-aspect retrieval strategy [Uncaptioned image] (detailed in Section 2.3). Finally, the retrieved data can optionally be assessed [Uncaptioned image] with novel metrics regarding how well it corresponds to the multi-aspect requirements (detailed in Section 3). As with the data preparation [Uncaptioned image] stage, the query execution [Uncaptioned image] stage is flexible, and the user can plug in their models [Uncaptioned image] / [Uncaptioned image] of choice and use their own queries [Uncaptioned image]. We also offer a dedicated synthetic query generator [Uncaptioned image] that can be used to construct multi-aspect input queries [Uncaptioned image] (detailed in Section 3) for evaluation purposes.
@@ 2.2 Constructing Multi-Aspect Embeddings [Uncaptioned image]
MRAG can leverage any embedding model with multi-head attention support to construct the multi-aspect embeddings for a given input text. In this work, we consider two embedding models from the MTEB leaderboard [16] as potential candidates. Specifically, the SFR-Embedding-Model [25] and the e5-mistral-7b-instruct [35], both based on the Mistral 7B architecture with 32 decoder blocks and 32 attention heads per multi-head attention.
@@ 2.3 Retrieval Strategies for Multi-Aspect Data [Uncaptioned image]
Algorithm 1 details the construction of importance scores. It is a heuristic based on extensive empirical evaluation; it gives high-quality results across the tested datasets and tasks. Intuitively, the score $s_{i}$ of a given head $h_{i}$ consists of two parts, $a_{i}$ and $b_{i}$. $a_{i}$ is the average of L2 norms of all embeddings in the vector space $i$; it represents how important a given head is: the larger the norms, the more attention was given to this attention head. $b_{i}$ is the average of cosine distances between all (or a randomly sampled subset, if the user wants to reduce pre-compute time) embeddings in vector space $i$. Intuitively, $b_{i}$ is a proxy for measuring the “spread” of vector space $i$: the larger $b_{i}$, the larger the average angle between different embeddings in this space is. Deriving $s_{i}$ as a product $a_{i}\cdot b_{i}$ ensures that we reward heads with high average attention and high average spread, but simultaneously penalize heads with lower average attention or with low average spread (both $a_{i}$ and $b_{i}$ are appropriately scaled).
@@
2.3.1 Integration with Data Stores
MRAG can be seamlessly used with different classes of data stores [Uncaptioned image] and nearest neighbor (NN) search approaches. It can be combined with both the exact and the approximate NN to find the matching (embedding, chunk)-pairs. These two parts of the broader RAG processing pipeline are orthogonal to MRAG.
@@ 3 Multi-Aspect Datasets, Queries, and Metrics
To assess how well MRAG performs on multi-aspect queries, and to compare it to modern RAG schemes, we need (1) datasets of documents that capture multi-aspectuality, (2) queries to the LLM that touch upon multi-aspectuality and require retrieving documents from the multi-aspect dataset, and (3) metrics that assess how well a RAG scheme retrieves such multi-aspect data. We now describe these three elements. In Section 4, we also briefly discuss real-world data and queries used.
Multi-Aspect Dataset Generation We first select conceptually different categories of documents. We primarily focus on publicly available Wikipedia articles and select 25 categories (e.g., countries, board games, historical swords, shipwrecks, etc.). For each category, we sample 50 documents. The first part of the document (overview) is used as a text chunk to be embedded. We enforce that each overview must have at least 800 characters, matching commonly used chunk sizes in RAG schemes.
@@
Multi-Aspect Query Generation We also require queries that touch upon a given number of $n$ aspects. For example, a query with 10 aspects must contain a question about 10 different documents from 10 different categories. We create such queries by selecting $n$ categories, sampling a document from each selected category (ensuring there are no duplicates overall), and then generating a story that combines these documents, using an LLM (GPT-3.5 Turbo). We construct 25 queries with 1, 5, 10, 15 and 20 aspects (125 queries in total). An example multi-aspect query sent to the LLM that requires retrieving 10 documents from 10 different categories, is pictured in the top part of Figure 3.
@@
Metrics We also design novel metrics to assess how well a given RAG scheme supports multi-aspectuality. For a query $Q$, a used retrieval strategy $S$ (detailed in Section 2.3), and $n$ documents from $n$ categories to retrieve, $Q_{rel}$ denotes the ideal set of documents that should be retrieved for $Q$. Then, $S(Q,n)$ is the set of the actually retrieved documents. We define the Retrieval Success Ratio as $\Xi(Q,n)=\frac{|S(Q,n)\cap Q_{rel}|}{|Q_{rel}|}$, i.e., the ratio of successfully retrieved relevant documents. Moreover, there is a case when a RAG scheme does not retrieve the exact desired document, but it still retrieves successfully some other document from the same category. To consider such cases, we use another measure, the Category Retrieval Success Ratio or $\Xi_{c}$. It has the same form as $\Xi(Q,n)$ above, with one difference: $S(Q,n)$ is now the set of all the retrieved documents that belong to categories of the ideal desired documents. Finally, to combine these two metrics, we use the Weighted Retrieval Success Ratio $\Xi_{w}$ as $\Xi_{w}=\frac{w\cdot\Xi+\Xi_{c}}{w+1}$. By varying $w$, the user can adjust the importance of exact document matches and category matches. An example of using these metrics to assess how well MRAG and Standard RAG capture multi-aspectuality is pictured in the bottom part of Figure 3.

### v2

@@ Abstract
Retrieval Augmented Generation (RAG) enhances the abilities of Large Language Models (LLMs) by enabling the retrieval of documents into the LLM context to provide more accurate and relevant responses. Existing RAG solutions do not focus on queries that may require fetching multiple documents with substantially different contents. Such queries occur frequently, but are challenging because the embeddings of these documents may be distant in the embedding space, making it hard to retrieve them all. This paper introduces Multi-Head RAG (MRAG), a novel scheme designed to address this gap with a simple yet powerful idea: leveraging activations of Transformer’s multi-head attention layer, instead of the decoder layer, as keys for fetching multi-aspect documents. The driving motivation is that different attention heads can learn to capture different data aspects. Harnessing the corresponding activations results in embeddings that represent various facets of data items and queries, improving the retrieval accuracy for complex queries. We provide an evaluation methodology and metrics, multi-aspect datasets that we release online, and real-world use cases to demonstrate MRAG’s effectiveness, showing improvements of up to 20% in relevance over standard RAG baselines. MRAG can be seamlessly integrated with existing RAG frameworks and benchmarking tools like RAGAS as well as different classes of data stores.
@@ 1 Introduction
Large Language Models (LLMs) transformed many machine learning tasks using in-context learning abilities. They achieved such accuracy by leveraging an increasing number of parameters, which in recent models have grown to hundreds of billions, making LLM training expensive in terms of both time and resources. It also comes with the danger of leaking confidential data into model weights (Yan et al., 2024; Shi et al., 2024; Patil et al., 2024). Additionally, continuous training through fine-tuning is necessary to keep LLMs up-to-date. Even using the newest data, LLMs display an ongoing problem of hallucinations (Zhang et al., 2023; Xu et al., 2024c; Huang et al., 2023) by providing factually incorrect information. Retrieval Augmented Generation (RAG) was proposed (Lewis et al., 2020; Guu et al., 2020) in order to address these issues as well as others and make LLMs more trustworthy.
The key idea behind RAG is to enhance the generative model’s capabilities by integrating a retrieval system that fetches relevant passages from a large corpus of data. In this setting, when a query is received, the retrieval system first identifies and retrieves pertinent information, which is fed into the generative model’s context for a more accurate and relevant response. Instead of the model storing information within its weights, RAG effectively leverages external knowledge, reducing hallucinations (by grounding the LLM reply in reliable sources), and ensuring that responses contain up-to-date knowledge (e.g., by accessing the Internet), all without requiring expensive training.
@@
Intense recent research efforts have been put into RAG (Gao et al., 2024; Zhao et al., 2024; Hu & Lu, 2024; Huang & Huang, 2024; Yu et al., 2024; Mialon et al., 2023; Li et al., 2022). On one hand, different RAG designs have been proposed, for example RAPTOR (Sarthi et al., 2024), Self-RAG (Asai et al., 2023), Chain-of-Note (Yu et al., 2023), and many others (Abdallah & Jatowt, 2024; Delile et al., 2024; Edge et al., 2024; Manathunga & Illangasekara, 2023; Zeng et al., 2024; Wewer et al., 2021; Xu et al., 2024b). In general, these schemes focus on making the retrieved data more accurate and relevant to the query. There have also been efforts into benchmarking and datasets for RAG evaluation (Chen et al., 2024b; Xiong et al., 2024; Lyu et al., 2024; Es et al., 2023).
@@
In this work, we propose Multi-Head RAG (MRAG): a scheme that addresses the above problem. Common practice in modern RAG designs is the use of embeddings based on last-layer decoder block activations. Our key idea is to use instead the activations of the multi-head attention part of the decoder block as embeddings. The Transformer architecture can be seen as a pipeline with many (e.g., 96 for GPT-3 (Wang et al., 2024b)) blocks, where a single block consists of an attention module and a feed-forward module. Each individual attention module is multi-headed: it consists of multiple parts called heads that learn different sets of weight matrices; see Figure 1 for an overview. It is conjectured that these different heads could capture different aspects of the processed data. We use this as a driving design feature that facilitates capturing the potential multi-aspectuality of the data without increasing space requirements compared to standard RAG, and without any fine-tuning or other modifications to the harnessed model (contribution 1).
Such multi-aspect embeddings are then directly used for both data items and query representation. Considering multi-aspectuality explicitly comes with challenges. For example, it is unclear how to assess whether a RAG solution does indeed harness multiple aspects when fetching documents. For this, we establish an evaluation methodology as well as a full data construction and query processing pipeline that implements the multi-aspect embedding idea (contribution 2). Our datasets facilitate broad evaluation by considering both fully-automatically generated, synthetic data as well as data based on specific industry use cases that show the benefits of MRAG (contribution 3). We ensure the relevance of our RAG datasets in real use cases by working directly with tech leaders (e.g., a generative AI division head) from 3 corporations, all of which actively use RAG in their own LLM infrastructures. Our evaluation illustrates the benefits in the relevance of retrieved documents, for example 20% over a modern RAG baseline for fetching multi-aspect Wikipedia articles, and comparable performance for single-aspect queries (contribution 4). We also show how MRAG and its benchmarking principles can be seamlessly integrated with both existing RAG solutions and benchmarking frameworks such as RAGAS (contribution 5). MRAG’s code is publicly available [^1] https://github.com/spcl/MRAG.
@@ 2 The MRAG Formulation & Pipeline
Decoder Formulation We first introduce formally the decoder architecture. We omit, for clarity, unnecessary details such as layer normalizations. The input is a text chunk that consists of $n$ tokens. The output of attention head $h$ for the $i$th token $x_{i}$ is defined as follows (Vaswani et al., 2017):
$\text{head}^{h}(\mathbf{x}_{i})=\sum_{j}w_{ij}\mathbf{v}_{j}^{h},\quad\text{where}$ (1)
$\vspace{-0.5em}w_{ij}=\text{softmax}\left(\left(\mathbf{q}_{i}^{h}\right)^{T}\mathbf{k}_{j}^{h}\right),\quad\mathbf{q}_{i}^{h}=\mathbf{W}_{q}^{h}\mathbf{x}_{i},\quad\mathbf{k}_{j}^{h}=\mathbf{W}_{k}^{h}\mathbf{x}_{j},\quad\mathbf{v}_{j}^{h}=\mathbf{W}_{v}^{h}\mathbf{x}_{j}$ (2)
where $\mathbf{W}_{q}^{h},\mathbf{W}_{k}^{h},\mathbf{W}_{v}^{h}$ are, respectively, learnable query, key, and value projections associated with head $h$, and $\mathbf{x}_{j}$ is the vector embedding of the $j$th token $x_{j}$. These outputs get combined to form the output of the $i$th multi-head attention block as follows:
@@
where matrix $\mathbf{W}_{o}$ is the linear layer that combines the outcomes of all the attention heads. This step is then followed by the Transformer feed-forward layer.
@@
Multi-Head RAG Formulation The key idea behind MRAG is simple: instead of the single activation vector generated by the last feed-forward decoder layer for the last token, we harness the $H$ separate activation vectors generated by the last attention layer for the last token, before processing it via $\mathbf{W_{o}}$. This can be formulated as a set of embeddings $\mathcal{S}$:
$\mathcal{S}=\{\mathbf{e}_{k}\forall_{k}\}\quad\text{where}\quad\mathbf{e}_{k}=\text{head}^{k}(\mathbf{x}_{n})$ (4)
which is simply the set of all outputs from the attention heads on the last token $\mathbf{x}_{n}$ of the input. As processing with multiple heads does not change the size of the output vector, $\mathcal{S}$ has the same space requirements as standard RAG. However, because we capture the separate embeddings before their mixing with $\mathbf{W}_{o}$, we conjecture that it gives more information about what the different parts of the input attend to, facilitating capturing multi-aspectuality.
@@ 2.1.1 Data Preparation
When preparing data [Uncaptioned image], we populate a data store [Uncaptioned image] with multi-aspect MRAG text embeddings [Uncaptioned image] and their corresponding documents [Uncaptioned image] or text chunks [Uncaptioned image] (MRAG is orthogonal to the type of data being embedded, and while we primarily use chunking of documents in order to reflect modern RAG pipelines, one can also embed whole documents or even other types of data). We create the multi-aspect embedding [Uncaptioned image] of each text chunk [Uncaptioned image] using a selected decoder-based embedding model [Uncaptioned image] (this part is detailed in Section 2.2). The user of the pipeline can plug in their model [Uncaptioned image] of choice as well as use their input data. We also offer a dedicated synthetic data generator [Uncaptioned image] that can be used to construct multi-aspect input documents [Uncaptioned image] (we detail this part in Section 3) for evaluation purposes.
@@ 2.1.2 Query Execution
During query execution [Uncaptioned image], we first generate a multi-aspect embedding [Uncaptioned image] of the input query [Uncaptioned image], using the selected embedding model [Uncaptioned image] (details in Section 2.2). Then, we find the nearest multi-aspect embeddings [Uncaptioned image] and their corresponding text chunks [Uncaptioned image] in the data store [Uncaptioned image] using a special multi-aspect retrieval strategy [Uncaptioned image] (detailed in Section 2.3). We ensure that there is no overhead in latency due to multiple aspects because computing these different smaller embeddings is done fully in parallel. Finally, the retrieved data can optionally be assessed [Uncaptioned image] with novel metrics regarding how well it corresponds to the multi-aspect requirements (detailed in Section 3). As with the data preparation [Uncaptioned image] stage, the query execution [Uncaptioned image] stage is flexible, and the user can plug in their models [Uncaptioned image] / [Uncaptioned image] of choice and use their own queries [Uncaptioned image]. We also offer a dedicated synthetic query generator [Uncaptioned image] that can be used to construct multi-aspect input queries [Uncaptioned image] (detailed in Section 3) for evaluation purposes.
@@ 2.2 Constructing Multi-Aspect Embeddings [Uncaptioned image]
MRAG can leverage any embedding model with multi-head attention support to construct the multi-aspect embeddings for a given input text. In this work, we consider two embedding models from the MTEB leaderboard (Huggingface, 2024) as potential candidates. Specifically, the SFR-Embedding-Model (Meng et al., 2024) and the e5-mistral-7b-instruct (Wang et al., 2024a), both based on the Mistral 7B architecture with 32 decoder blocks and 32 attention heads per multi-head attention.
@@ 2.3 Retrieval Strategies for Multi-Aspect Data [Uncaptioned image]
Algorithm 1 details the construction of importance scores. It is a heuristic based on extensive empirical evaluation; it gives high-quality results across the tested datasets and tasks. Intuitively, the score $s_{i}$ of a given head $h_{i}$ consists of two parts, $a_{i}$ and $b_{i}$. $a_{i}$ is the average of L2 norms of all embeddings in the vector space $i$; it represents how important a given head is: the larger the norms, the more attention was given to this attention head. $b_{i}$ is the average of cosine distances between all (or a randomly sampled subset, if the user wants to reduce pre-compute time) embeddings in vector space $i$.
@@
Intuitively, $b_{i}$ is a proxy for measuring the “spread” of vector space $i$: the larger $b_{i}$, the larger the average angle between different embeddings in this space is. Deriving $s_{i}$ as a product $a_{i}\cdot b_{i}$ ensures that we reward heads with high average attention and high average spread, but simultaneously penalize heads with lower average attention or with low average spread (both $a_{i}$ and $b_{i}$ are appropriately scaled).
@@
2.3.1 Integration with Data Stores & Other Types of Models
MRAG can be seamlessly used with different classes of data stores [Uncaptioned image] and nearest neighbor (NN) search approaches. It can be combined with both the exact and the approximate NN to find the matching (embedding, chunk)-pairs. These two parts of the broader RAG processing pipeline are orthogonal to MRAG. Similarly, MRAG does not depend specifically on the embedding form, as long as it is based on a model that harnesses multi-head attention any such approach that results in a valid embedding can be used. As such, it could also be used with models such as RetroMAE (Xiao et al., 2022) and the classic BGE-embeddings (Xiao et al., 2022; Chen et al., 2024a).
@@ 3 Multi-Aspect Datasets, Queries, and Metrics
To assess how well MRAG performs on multi-aspect queries, and to compare it to modern RAG schemes, we need (1) datasets of documents that capture multi-aspectuality, (2) queries to the LLM that touch upon multi-aspectuality and require retrieving different documents from the multi-aspect dataset, and (3) metrics that asses how well a given RAG scheme retrieves such multi-aspect data. We now describe these three elements. In Section 4, we also discuss real-world data and queries used to evaluate MRAG.
3.1 Multi-Aspect Datasets
We first select conceptually different categories of documents for a synthetic dataset. Here, we harness publicly available Wikipedia articles. In the dataset construction pipeline, the user selects a given number of categories (e.g., countries, board games, historical swords, shipwrecks, etc.) and then, for each category, they sample a specified number of documents. The first part of the document (overview) is used as a text chunk to be embedded. We enforce that each overview must have at least 800 characters, matching commonly used chunk sizes in RAG schemes. We also use multi-aspect real-world inspired datasets consisting of NDAs and reports describing industry accidents in chemical processing plants. We ensure the usefulness of these datasets by working directly with tech leaders from 3 corporations that rely on RAG in their in-house LLM-driven report generation and analytics frameworks. Example categories of the legal documents are legal areas (energy law, family law, criminal law, etc.) or document language style (aggressive, mild, neutral, etc.). Examples of accident causes are natural disasters, human mistakes, or lack of proper training. We fully release these datasets to propel RAG research. Details on all three datasets can be found in the Appendix B.2. In our evaluation, we use a total of 13,750 documents.
3.2 Multi-Aspect Query Generation
@@
We also require queries that touch upon a given number of $n$ aspects. For example, a query with 10 aspects must contain a question about 10 different documents from 10 different categories. We create such queries by selecting $n$ categories, sampling a document from each selected category (ensuring there are no duplicates overall), and then generating a story that combines these documents, using an LLM (GPT-3.5 Turbo). We construct 25 queries with 1, 5, 10, 15 and 20 aspects (125 queries in total). An example multi-aspect query sent to the LLM that requires retrieving 10 documents from 10 different categories, is pictured in the top part of Figure 3.
3.3 Metrics
@@
We also design novel metrics to assess how well a given RAG scheme supports multi-aspectuality. For a query $Q$, a used retrieval strategy $S$ (detailed in Section 2.3), and $n$ documents from $n$ categories to retrieve, $Q_{rel}$ denotes the ideal set of documents that should be retrieved for $Q$. Then, $S(Q,n)$ is the set of the actually retrieved documents. We define the Retrieval Success Ratio as $\Xi(Q,n)=\frac{|S(Q,n)\cap Q_{rel}|}{|Q_{rel}|}$, i.e., the ratio of successfully retrieved relevant documents. Moreover, there is a case when a RAG scheme does not retrieve the exact desired document, but it still retrieves successfully some other document from the same category. While less desired, it still increases chances for a more accurate LLM answer following the retrieval. For example, when asking the LLM to determine the cause of an industry accident, fetching the documents in the same category as the accident being queried about, improves the chances for the LLM to give a more relevant answer. To consider such cases, we use another measure, the Category Retrieval Success Ratio or $\Xi_{c}$. It has the same form as $\Xi(Q,n)$ above, with one difference: $S(Q,n)$ is now the set of all the retrieved documents that belong to categories of the ideal desired documents. Finally, to combine these two metrics, we use the Weighted Retrieval Success Ratio $\Xi_{w}$ as $\Xi_{w}=\frac{w\cdot\Xi+\Xi_{c}}{w+1}$. By varying $w$, the user can adjust the importance of exact document matches and category matches. An example of using these metrics to assess how well MRAG and Standard RAG capture multi-aspectuality is pictured in the bottom part of Figure 3.